OpenStudy (rebecca1588):

What is the solution to the system of equations? y=2x+6 3x-y=6 ( ? , ? ) <<<<

OpenStudy (anonymous):

That is correct, Rebecca.

OpenStudy (cathyangs):

They give you y in terms of x already. Plug that into the second equation, to find an equation with just x and numbers. I get 3x - (2x + 6) = 6 x -6 = 6 x = 12 plugging that back into the 1st equation, y = 2*12 + 6 = 30. You're right!

OpenStudy (rebecca1588):

Thanks :D